OpenStudy (anonymous):

Medal and fan if you can give me the answer!!!!!!!!!!!!!!!!!One Solution To The Problem Below Is 7. What Is The Other Solution? Y^2-49=0

OpenStudy (sweetburger):

This isnt too tough if you remember that \[\sqrt{4}= +2 or -2\]

OpenStudy (sweetburger):

(-2)^2=4 and (2)^2 = 4 make sense?

OpenStudy (sweetburger):

@Lyssa1500 do you have an idea what the answer is?

OpenStudy (anonymous):

I think so hang on

OpenStudy (anonymous):

ok is that all I do?

OpenStudy (sweetburger):

what did you get?

OpenStudy (anonymous):

@Lyssa1500 answer him back